OpenStudy (anonymous):

express square root of 50-the square root of 18 + the square root of 8

OpenStudy (solomonzelman):

you have to convert this word phrase into an equation?

OpenStudy (anonymous):

\[\sqrt{50}-\sqrt{18}+\sqrt{8} ?\]

OpenStudy (anonymous):

yes